If a bullet leaves the muzzle of a rifle with a speed of 600m/s, and the barrel ofthe rifle is 0.9m long, what is the acceleration of the bullet while in the barrel?
Answer #1
GivenThe initial velocity of the bullet vo= 0Themuzzle of a rifle with a speed v =600m/sDistance x= 0.90 m---------------------------------------------------------------------From the kinematic relations,v2 =vo2+ 2axThe acceleration of the bullet while in the barrel isa = [v2-vo2]/(2x)= [(600 m/s)2 -0]/(2(0.90m))=2.0x105m/s2
answered by: Kharato
Answer #2

can you guy show me the work